# Shadow of Quantum Improved Regular Kerr Black Hole and parameter constrains with EHT observations

### v1

@@ Abstract
Quantum Improved Regular Kerr (QIRK) Black Hole is a rotating regular black hole based on the asymptotic safety method. This black hole not only resolves ring singularity and avoids closed timelike curves, but also has well defined thermodynamics. Therefore, it is crucial to find some observable features of this rotating black hole. In this article, we numerically determine the specific parameter range of the QIRK black hole after ensuring the three key properties mentioned above, while investigating its black hole shadow, we find that the extremal QIRK black hole, under a critical angular momentum $a_{\mathrm{cri}}$, can have a similar shadow to the non extremal Kerr black hole. Furthermore, with recent observations from the Event Horizon Telescope (EHT) of Sgr A* and earlier observations of the supermassive black hole M87*, we constrain the QIRK black hole using observational data and explore its potential as an astronomical object.
@@ I Introduction
Black holes, predicted by general relativity, are one of the most enigmatic objects in the universe. Classical black holes, such as the Kerr black hole kerr_gravitational_1963 which describes a rotating black hole, play a crucial role in understanding numerous astrophysical phenomena, including quasars and gamma-ray bursts. However, a significant issue with these classical solutions is the presence of singularities, characterized by the divergence of curvature invariants and geodesic incompleteness wald_general_1984 ; hawking_large_1973 . To address this problem, regular black holes (RBHs) were introduced. Unlike traditional black holes, they are free from essential singularities throughout the entire spacetime bardeen_non-singular_1968 ; dymnikova_vacuum_1992 ; ayon-beato_regular_1998 ; bronnikov_regular_2001 . The first regular black hole model was proposed by Bardeen bardeen_non-singular_1968 , and was later interpreted as a solution derived from Einstein’s field equations with nonlinear electromagnetic field Ayon-Beato:2000mjt . This approach has been extended to explain various spherically symmetric RBH models Ayon-Beato:2000mjt ; PhysRevLett.96.031103 ; bronnikov_regular_2001 ; fan_construction_2016 ; bronnikov_regular_2006 ; bronnikov_field_2022 . Since Bardeen’s work, numerous other regular black hole models have been developed Ayon-Beato:1999kuh ; burinskii_new_2002 ; dymnikova_regular_2004 ; berej_regular_2006 ; junior_regular_2015 ; sajadi_nonlinear_2017 ; toshmatov_comment_2018 ; bambi_rotating_2013 ; toshmatov_rotating_2014 ; azreg-ainou_generating_2014 ; ghosh_nonsingular_2015 ; ghosh_radiating_2015 , including several rotating RBH models bambi_rotating_2013 ; toshmatov_rotating_2014 ; azreg-ainou_generating_2014 ; ghosh_nonsingular_2015 ; ghosh_radiating_2015 .
@@
Two primary methods have been established for constructing regular black hole models. The first one involves solving Einstein’s field equations with specific matter sources noauthor_noncommutative_2006 ; noauthor_non-commutative_2009 ; noauthor_non-minimal_2007 . The second method derives RBHs as quantum corrections to classical black holes with singularities. The notable examples come from frameworks such as loop quantum gravitymodesto_disappearance_2004 ; gambini_black_2008 ; perez_black_2017 . A promising alternative is the asymptotic safety scenario Reuter:1996cp ; Souma:1999at ; doi:10.1142/10369 ; 2019qgfr.book…..R ; bonanno_renormalization_2000 ; koch_black_2014 ; chen2024quantumimprovedregularkerr , which suggests that quantum gravity effects, particularly through the functional renormalization group, introduce a repulsive force near black hole cores bonanno_renormalization_2000 ; Bonanno:2006eu , potentially resolving singularities. Many such approaches have been studied for various black hole solutions reuter_running_2004 ; falls_black_2012 ; platania_black_2023 ; eichhorn_black_2022 ; koch_structural_2013 ; pawlowski_quantum-improved_2018 ; ishibashi_quantum_2021 ; Litim:2013gga ; Harst:2011zx ; Falls:2012nd ; Adeifeoba:2018ydh ; 10.3389/fphy.2020.00188 ; Borissova:2022mgd ; Ruiz:2021qfp .
@@
The observation of light deflection in a gravitational field in 1919 provided the first experimental confirmation of a prediction from general relativity. This phenomenon, now known as gravitational lensing, is a key method for detecting weak gravitational fields. When a photon passes near a black hole, it can either be absorbed, bounded in its vicinity, or scattered to infinity. The observed image of such photons is referred to as the black hole shadow. The first analytic study of black hole shadows began with Synge’s Synge:1966okc and Luminet’s Luminet:1979nyg discussions of the Schwarzschild black hole, followed by Bardeen’s analysis of Kerr black hole in 1973 1973blho.conf..215B . These foundational works paved the way for further exploration of black hole shadows in various models noauthor_escape_nodate ; grenzebach_photon_2014 ; bambi_apparent_2009 ; takahashi_shapes_2004 ; wei_observing_2013 ; abdujabbarov_shadow_2013 ; atamurotov_optical_2015 ; schee_optical_2009 ; tsukamoto_constraining_2014 ; takahashi_black_2005 ; bambi_direct_2012 ; khodadi_black_2020 ; shaikh_shadows_2019 ; falcke_toward_2013 ; tsupko_analytical_2017 ; qian_cuspy_2022 ; noauthor_charged_2022 ; contreras_black_2020 ; perlick_calculating_2022 ; falcke_viewing_1999 ; atamurotov_shadow_2013 ; vries_apparent_2000 ; hioki_measurement_2009 ; abdujabbarov_coordinate-independent_2015 ; yumoto_shadows_2012 ; Cunha:2015yba ; Cunha:2018acu ; Lima:2021las ; Chen:2020aix ; 2005Natur.438…62S ; Cao:2024kht , including regular black holes li_measuring_2014 ; amir_shapes_2016 ; kumar_shadow_2019 ; banerjee_signatures_2022 ; allahyari_magnetically_2020 ; olmo_shadows_2023 ; abdujabbarov_shadow_2016 ; tsukamoto_black_2018 ; dymnikova_identification_2019 ; noauthor_ergosphere_2020 ; jusufi_quasinormal_2021 ; ling_shadows_2022 ; Kumar:2018ple ; Kumar:2020yem ; Kumar:2020ltt ; Cao:2023par .
@@
A major breakthrough in the study of black hole shadows came in 2019 EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ggy ; EventHorizonTelescope:2019pgp ; EventHorizonTelescope:2019ths ; EventHorizonTelescope:2019uob ; EventHorizonTelescope:2019jan when the Event Horizon Telescope (EHT) released the first image of the supermassive black hole M87*. This was followed in 2022 by EHT’s observations of Sgr A* EventHorizonTelescope:2022wkp ; EventHorizonTelescope:2022xqj ; EventHorizonTelescope:2022urf ; EventHorizonTelescope:2022wok ; EventHorizonTelescope:2022apq ; EventHorizonTelescope:2022exc . Based on these EHT results, the shadows of M87* and Sgr A* have become important tools for testing and constraining theories of gravity in the strong-field regime. Recently, a vast body of literature aimed at deepening our understanding of black holes through these observations afrin_parameter_2021 ; vagnozzi_hunting_2019 ; afrin_estimating_2022 ; vagnozzi_horizon-scale_2023 ; kumar_rotating_2020 ; kumar_gravitational_2020 ; walia_testing_2022 ; ghosh_parameters_2021 ; islam_investigating_2023 ; noauthor_shadows_nodate ; walia_observational_2023 ; Cunha:2019ikd ; Kumar:2020yem ; Kumar:2020ltt ; Uniyal:2022vdu ; Pantig:2022ely ; Afrin:2021wlj ; Bambi:2019tjh .
Given the importance of black hole shadows and the recent EHT observations, this article investigates the QIRK black hole. By hypothesizing that real black holes may conform to this model, we explore its expected observational signatures. Our research highlights the novel properties of the QIRK black hole’s shadow, particularly under extremal conditions, where it significantly differs from that of an extremal Kerr black hole. Furthermore, we use EHT data to constrain the parameters of QIRK black holes, aiming for a more comprehensive and detailed understanding of this model.
This article is organized as follows: Section II examines the constraints imposed by QIRK black holes, including their horizons, regularity, and avoidance of closed timelike curves. Section III discusses the motion of light in both the radial and angular directions. Section IV introduces the shadow correlation function, which facilitates the discussion of shadows in Section V. Section VI analyzes black hole shadows by using EHT data to further constrain the parameters of QIRK black holes. Finally, Section VII summarizes the article and discusses the potential candidature of QIRK black holes, through shadow observations, as viable models for astrophysical black holes.
@@
The asymptotically safe scenario for a quantum generalization of general relativity with a cosmological constant proposes an energy-scale $k$ dependent Newton constant $G(k)$ and cosmological constant $\Lambda(k)$. Assuming the cosmological constant is already at its fixed point and negligibly small, the solution for the Newtonian coupling derived from the associated renormalization group equations takes the form pawlowski_quantum-improved_2018 :
@@
where $\omega$ is a constant of order 1 representing quantum effects, and $G_{0}=G(0)$ is the Newton constant.
Based on resolving singularities and ensuring the consistency of black hole thermodynamics, Ref. chen2024quantumimprovedregularkerr also incorporates some physical discussions, namely that quantum corrections in graviy theory should become significant at the Planck scale. They adopted the following identification:
@@
Here, $G_{0}$ is the Newton constant, $M$ is the mass parameter, $a$ is the rotation parameter, and $\widetilde{\omega}$ is the quantum correction parameter introduced by the QIRK metric. The magnitude of $\widetilde{\omega}$ reflects the strength of the quantum effects, and we require $\widetilde{\omega}\geq 0$. When $\widetilde{\omega}=0$, the solution reduces to the classical Kerr case.
For ease of discussion, we will introduce dimensionless physical quantities as follows:
@@
In the following discussion, we will use the dimensionless variables $x$, $A$, and $\widetilde{\Omega}$. However, for convenience, we will still denote them by the original symbols $r$, $a$, and $\widetilde{\omega}$.
In this section, we will discuss the specific conditions required to resolve the singularities, ensure the existence of horizons and avoid the closed timelike curves. We will examine the implications of each condition and their corresponding physical significance. Finally, these elements will be integrated to provide a comprehensive understanding of the QIRK metric.
@@ II.1 Resolving the Singularity and Ensuring Horizon Existence
Figure 1: we’ve chosed $a=0.9$ and $p=4$ to analysize the effect of $\widetilde{\omega}$ on the behavior of $\Delta(r)$. $\widetilde{\omega}_{c}$ is the critical vaule which two horizons will coincident. If $\widetilde{\omega}>\widetilde{\omega}_{c}$, then there are no horizons. On the contrary, there exists at least one horizon.
@@ II.2 Avoiding Closed Timelike Curves
It is well known that closed timelike curves (CTCs) exist in the Kerr black hole, which could lead to violations of causality. However, the CTCs are eliminated in this black hole model due to quantum effects. In this section, we examine the parameter constraints that arise from the avoidance of CTCs.
@@
Firstly, considering that the $g_{\varphi\varphi}$ component contains the $\Delta(r)$ term, even when $\theta=\pi/2$, its expression remains challenging to simplify due to the higher powers of $G(r)$ in $\Delta(r)$. Secondly, to accurately determine the parameter ranges that allow us to utilize the Event Horizon Telescope (EHT) data for tighter parameter constraints in Sec. VI, we adopt a numerical approach to analyze the conditions required to avoid CTCs.
@@
$g_{\varphi\varphi}=\frac{((r^{2}+a^{2})^{2}-a^{2}\Delta\sin^{2}\theta)\sin^{2}\theta}{\Sigma},\qquad\Sigma=r^{2}+a^{2}\cos^{2}\theta.$ (11)
@@
Figure 2: Plot of $\psi(r)$ as a function of $r$ for $a=0.9$ and $p=4$. When $\widetilde{\omega}>\widetilde{\omega}_{m}$, $\psi(r)$ has no zero points, corresponding to $g_{\varphi\varphi}>0$ for all $r$. Conversely, if $\psi(r)$ has a zero point, there will exist some $r$ for which $g_{\varphi\varphi}<0$, implying the existence of CTCs.
@@
Additionally, the green-shaded area denotes the region where no horizon exists, but CTCs are present. This suggests that certain causality-violating events or signals may be observed by external observers. While such a region is permitted by the QIRK metric, its interpretation in the real world raises significant questions and therefore warrants further investigation.
@@
In this paper, we focus on the cases of $p=4$ and $p=6$ for the following reasons: first, as illustrated in Fig. 3, as the value of $p$ increases, the parameter range for the QIRK black hole expands, and the orange curve representing the avoidance of CTCs approaches the $a$-axis, with $a_{\text{max}}$ gradually approaching 1. This could potentially obscure certain distinctive features of the QIRK black hole. Second, from the aspect of black hole shadow, which we will discuss in Sec. V, a smaller value of $p$ results in greater deviations from the Kerr black hole. Thus, considering smaller values of $p$ better highlights the differences between the QIRK black hole and the Kerr black hole.
@@
Table 1: Maximum rotation parameter $a_{\text{max}}$ and critical rotation parameter $a_{\text{cri}}$ for different values of $p$.
@@
$a_{\text{cri}}$ 0.873588 0.921738 0.9466 0.961165 0.970453 0.976748 0.981217
III Null geodesics and Photon sphere
When photons move around a black hole, the strong gravitational field causes them to bend. Among the many possible orbits, the most unique is the spherical photon orbit, which defines the boundary between the shaded and the illuminated region. Therefore, studying the motion of photon is crucial for analyzing the structure and observational properties of the black hole shadow.
Before proceeding with the discussion, we would like to clarify the following: we have observed that for many rotating black holes, their null geodesic equations, including the photon sphere, share a similar form to that of the Kerr black hole, with the only difference being a correction to $\Delta$. Considering that, first, we have already discussed the behavior of $\Delta$ for the QIRK black hole in Fig. 1, and second, substituting the specific form of $\Delta$ for the QIRK black hole would make the expressions unnecessarily lengthy and less intuitive, we will not include the explicit form of $\Delta$ in this section. This approach simplifies the generalization to other rotating black holes. Moreover, the analytical expressions derived from the analysis of the photon sphere structure in this section serve as a foundation for intuitively identifying the photon sphere in the next section.

### v2

# Image of Quantum Improved Regular Kerr Black Hole and Parameter Constraints from EHT Observations

@@ Abstract
Quantum Improved Regular Kerr (QIRK) black hole is a rotating regular black hole model constructed based on the asymptotic safety method. The model eliminates the ring singularity and prevents the formation of closed timelike curves, while retaining well-defined thermodynamic properties. Given these properties, probing the observable features of the QIRK black hole is important. In this work, we numerically determine the region of parameter space in which the QIRK spacetime remains regular, admits an event horizon, and is free of closed timelike curves. Subsequently, we simulate images of a QIRK black hole surrounded by a thin accretion disk. We find the primary effect of the quantum correction parameter, $\widetilde{\omega}$, is a systematic reduction in the overall observed intensity, with only subtle effects on the image geometry. Using observational data from the Event Horizon Telescope (EHT) for Sgr A* and M87*, we further constrain the parameters of the QIRK black hole. Moreover, since there exist QIRK parameters that are free of singularities and can admit closed timelike curves, we investigate the images of CTCs under these conditions. These results reveal the distinctive observational features of the QIRK spacetime and provide a quantitative basis for assessing its viability as an astrophysical candidate.
@@ I Introduction
Black holes, predicted by general relativity, are one of the most enigmatic objects in the universe. Classical black holes, such as the Kerr black hole kerr_gravitational_1963 which describes a rotating black hole, play a crucial role in understanding numerous astrophysical phenomena, including quasars and gamma-ray bursts. However, a significant issue with these classical solutions is the presence of singularities, characterized by the divergence of curvature invariants and geodesic incompleteness wald_general_1984 ; hawking_large_1973 . To address this problem, regular black holes (RBHs) were introduced. Unlike traditional black holes, they are free from essential singularities throughout the entire spacetime bardeen_non-singular_1968 ; dymnikova_vacuum_1992 ; ayon-beato_regular_1998 ; bronnikov_regular_2001 . The first regular black hole model was proposed by Bardeen bardeen_non-singular_1968 , and was later interpreted as a solution derived from Einstein’s field equations with nonlinear electromagnetic field Ayon-Beato:2000mjt . This approach has been extended to describe various spherically symmetric RBH models Ayon-Beato:2000mjt ; PhysRevLett.96.031103 ; bronnikov_regular_2001 ; fan_construction_2016 ; bronnikov_regular_2006 ; bronnikov_field_2022 . Since Bardeen’s work, numerous other regular black hole models have been developed Ayon-Beato:1999kuh ; burinskii_new_2002 ; dymnikova_regular_2004 ; berej_regular_2006 ; junior_regular_2015 ; sajadi_nonlinear_2017 ; toshmatov_comment_2018 ; bambi_rotating_2013 ; toshmatov_rotating_2014 ; azreg-ainou_generating_2014 ; ghosh_nonsingular_2015 ; ghosh_radiating_2015 , including several rotating RBH models bambi_rotating_2013 ; toshmatov_rotating_2014 ; azreg-ainou_generating_2014 ; ghosh_nonsingular_2015 ; ghosh_radiating_2015 .
@@
Two primary methods have been established for constructing regular black hole models. The first one involves solving Einstein’s field equations with specific matter sources noauthor_noncommutative_2006 ; noauthor_non-commutative_2009 ; noauthor_non-minimal_2007 . The second method derives RBHs as quantum corrections to classical black holes with singularities. The notable examples come from frameworks such as loop quantum gravitymodesto_disappearance_2004 ; gambini_black_2008 ; perez_black_2017 ; Battista:2023iyu . A promising alternative is the asymptotic safety scenario Reuter:1996cp ; Souma:1999at ; doi:10.1142/10369 ; 2019qgfr.book…..R ; bonanno_renormalization_2000 ; koch_black_2014 ; chen2024quantumimprovedregularkerr , which suggests that quantum gravity effects, particularly through the functional renormalization group, introduce a repulsive force near black hole cores bonanno_renormalization_2000 ; Bonanno:2006eu , potentially resolving singularities. Many such approaches have been studied for various black hole solutions reuter_running_2004 ; falls_black_2012 ; platania_black_2023 ; eichhorn_black_2022 ; koch_structural_2013 ; pawlowski_quantum-improved_2018 ; ishibashi_quantum_2021 ; Litim:2013gga ; Harst:2011zx ; Falls:2012nd ; Adeifeoba:2018ydh ; 10.3389/fphy.2020.00188 ; Borissova:2022mgd ; Ruiz:2021qfp .
@@
The observation of light deflection in a gravitational field in 1919 provided the first experimental confirmation of a prediction from general relativity. This phenomenon, now known as gravitational lensing Ghosh:2022mka ; AbhishekChowdhuri:2023ekr , is a key method for detecting gravitational fields. When a photon passes near a black hole, it can either be absorbed, trapped in its vicinity, or scattered to infinity. The observed image of such photons is referred to as the black hole shadow  [^5] Here, ”shadow” specifically refers to the photon ring image of the black hole, not the inner shadow produced by a thin accretion disk. . The first analytic study of black hole shadows began with Synge’s Synge:1966okc and Luminet’s Luminet:1979nyg discussions on the Schwarzschild black hole, followed by Bardeen’s analysis of the Kerr black hole in 1973 1973blho.conf..215B . These foundational works paved the way for further exploration of black hole shadows in various models noauthor_escape_nodate ; grenzebach_photon_2014 ; bambi_apparent_2009 ; takahashi_shapes_2004 ; wei_observing_2013 ; abdujabbarov_shadow_2013 ; atamurotov_optical_2015 ; schee_optical_2009 ; tsukamoto_constraining_2014 ; takahashi_black_2005 ; bambi_direct_2012 ; khodadi_black_2020 ; shaikh_shadows_2019 ; falcke_toward_2013 ; tsupko_analytical_2017 ; qian_cuspy_2022 ; noauthor_charged_2022 ; contreras_black_2020 ; perlick_calculating_2022 ; falcke_viewing_1999 ; atamurotov_shadow_2013 ; vries_apparent_2000 ; hioki_measurement_2009 ; abdujabbarov_coordinate-independent_2015 ; yumoto_shadows_2012 ; Cunha:2015yba ; Cunha:2018acu ; Lima:2021las ; Chen:2020aix ; 2005Natur.438…62S ; Cao:2024kht ; Chowdhuri:2020ipb ; Konoplya:2024lch , including regular black holes li_measuring_2014 ; amir_shapes_2016 ; kumar_shadow_2019 ; banerjee_signatures_2022 ; allahyari_magnetically_2020 ; olmo_shadows_2023 ; abdujabbarov_shadow_2016 ; tsukamoto_black_2018 ; dymnikova_identification_2019 ; noauthor_ergosphere_2020 ; jusufi_quasinormal_2021 ; ling_shadows_2022 ; Kumar:2018ple ; Kumar:2020yem ; Kumar:2020ltt ; Cao:2023par ; Ghosh:2022gka .
Although the black hole shadow offers a direct probe of strong gravitational fields, observed black hole images depend not only on spacetime geometry but also on emission from accreting material and gravitational lensing. The thin accretion disk model, with its clear physical assumptions and sensitivity to lensing effects, has become an important framework for studying black hole imaging. In this model, photons emitted from different regions of the disk follow geodesics to the observer, producing features such as a direct image of the disk, higher-order lensing rings, and the narrow photon ring. These features encode both the black hole’s spacetime geometry and the emission properties. By studying black hole images under a thin accretion disk, one can separate lensing induced features such as the inner shadow (horizon image) and the photon ring, thereby measuring black hole parameters, and testing possible deviations from general relativity Chael:2021rjo ; Hou:2022eev ; Meng:2025ivb . In particular, for quantum-corrected models like QIRK, thin-disk imaging may reveal differences from Kerr black holes, providing a new perspective for seeking signatures of quantum gravity effects.
@@
A major breakthrough in the study of black hole images came in 2019 EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ggy ; EventHorizonTelescope:2019pgp ; EventHorizonTelescope:2019ths ; EventHorizonTelescope:2019uob ; EventHorizonTelescope:2019jan when the Event Horizon Telescope (EHT) released the first image of the supermassive black hole M87*. This was followed in 2022 by EHT’s observations of Sgr A* EventHorizonTelescope:2022wkp ; EventHorizonTelescope:2022xqj ; EventHorizonTelescope:2022urf ; EventHorizonTelescope:2022wok ; EventHorizonTelescope:2022apq ; EventHorizonTelescope:2022exc . Based on these EHT results, the images of M87* and Sgr A* have become important tools for testing and constraining theories of gravity in the strong-field regime. Recently, a vast body of literature has aimed at deepening our understanding of black holes through these observations afrin_parameter_2021 ; vagnozzi_hunting_2019 ; afrin_estimating_2022 ; vagnozzi_horizon-scale_2023 ; kumar_rotating_2020 ; kumar_gravitational_2020 ; walia_testing_2022 ; ghosh_parameters_2021 ; islam_investigating_2023 ; noauthor_shadows_nodate ; walia_observational_2023 ; Cunha:2019ikd ; Kumar:2020yem ; Kumar:2020ltt ; Uniyal:2022vdu ; Pantig:2022ely ; Afrin:2021wlj ; Bambi:2019tjh ; Shaikh:2021yux ; Pal:2023wqg ; Aliyan:2024xwl ; Nozari:2024jiz ; Nozari:2023flq .
Motivated by these developments, we first simulate the images of QIRK black holes illuminated by a thin accretion disk, to quantify how the quantum correction parameter $\widetilde{\omega}$ affects black hole images. Furthermore, utilizing EHT observational data for M87* and Sgr A*, we impose constraints on the QIRK black hole parameters (spin $a$ and quantum correction $\widetilde{\omega}$), to explore their possible range of existence in the real universe. Finally, we investigate the optical characteristics of closed timelike curves in a special, horizonless QIRK spacetime, to better understand how quantum corrections modify the interior ($r\!\leq\!0$) geometry.
This article is organized as follows: Section II establishes the framework of the QIRK, discussing the conditions for regularity, the existence of horizons, and the requirements to avoid closed timelike curves (CTCs) to define the physically relevant parameter space. Section III analyses null geodesics and the photon-sphere structure. Section IV provides a detailed analysis of QIRK black hole images illuminated by a thin accretion disk. Section V uses Event Horizon Telescope (EHT) observations of M87* and Sgr A* to impose constraints on the QIRK black hole parameters. Section VI explores the optical features of CTCs in horizonless QIRK spacetimes. Finally, Section VII summarizes the main findings and discusses the potential of the QIRK black hole as an astrophysical black hole candidate under current and future observational capabilities.
@@
The asymptotically safe scenario for a quantum generalization of general relativity with a cosmological constant proposes an energy-scale $k$ dependent Newton constant $G(k)$ and a cosmological constant $\Lambda(k)$. Assuming the cosmological constant is already at its fixed point and negligibly small, the solution for the Newtonian coupling derived from the associated renormalization group equations takes the form pawlowski_quantum-improved_2018 :
@@
where $\omega$ is a constant of order 1 representing quantum effects, and $G_{0}=G(0)$ is the usual Newton constant.
Besides resolving singularities and ensuring the consistency of black hole thermodynamics, Ref. chen2024quantumimprovedregularkerr additionally shows that quantum corrections dominate near the Planck scale, motivating the identification:
@@
Here, $G_{0}$ is the Newton constant as before, $M$ is the mass parameter, $a$ is the rotation parameter, and $\widetilde{\omega}$ is the quantum correction parameter introduced below Eq.(2). The magnitude of $\widetilde{\omega}$ reflects the strength of the quantum effects, and we require $\widetilde{\omega}\geq 0$. Obviously, the solution reduces to the case of Kerr when $\widetilde{\omega}=0$.
To simplify the discussion, one can introduce three dimensionless physical quantities as follows:
@@
In the following discussion, these dimensionless variables $x$, $A$, and $\widetilde{\Omega}$ will be substantially used. However, for convenience, we will still denote them by the original symbols $r$, $a$, and $\widetilde{\omega}$.
In this section, we discuss the conditions necessary to remove singularities, guarantee the presence of horizons, and prevent closed timelike curves. We then examine the implications of each condition and their physical significance. Finally, we integrate these results to present a comprehensive understanding of the QIRK metric.
@@ II.1 Resolving the Singularity and Ensuring Horizon Existence
Figure 1: We’ve chosen $a=0.9$ and $p=4$ to analyze the effect of $\widetilde{\omega}$ on the behavior of $\Delta(r)$. $\widetilde{\omega}_{c}$ is the critical value which two horizons will coincident. If $\widetilde{\omega}>\widetilde{\omega}_{c}$, then there are no horizons. On the contrary, there exists at least one horizon.
@@ II.2 Avoiding Closed Timelike Curves
It is well known that closed timelike curves (CTCs) exist in the Kerr black hole, which could lead to violations of causality. However, the CTCs can be eliminated in this black hole model due to quantum effects. In this section, we examine the parameter constraints that arise from the avoidance of CTCs.
@@
Firstly, considering that the $g_{\varphi\varphi}$ component contains the $\Delta(r)$ term, even when $\theta=\pi/2$, its expression remains challenging to simplify due to the higher powers of $G(r)$ in $\Delta(r)$. Secondly, to accurately determine the parameter ranges that allow us to utilize the Event Horizon Telescope (EHT) data for tighter parameter constraints in Sec. V, we adopt a numerical approach to analyze the conditions required to avoid CTCs.
@@
$g_{\varphi\varphi}=\frac{((r^{2}+a^{2})^{2}-a^{2}\Delta\sin^{2}\theta)\sin^{2}\theta}{\Sigma}.$ (11)
@@
Figure 2: Plot of $\psi(r)$ as a function of $r$ for $a=0.9$ and $p=4$. When $\widetilde{\omega}>\widetilde{\omega}_{m}$, $\psi(r)$ has no real zeros, corresponding to $g_{\varphi\varphi}>0$ for all $r$. Conversely, if $\psi(r)$ has a zero point, there will exist some $r$ for which $g_{\varphi\varphi}<0$, implying the existence of CTCs.
@@
Additionally, the green-shaded area denotes the region where no horizon exists, but CTCs are present. This suggests that certain causality-violating events or signals may be observed by external observers. While such a region is allowed by the QIRK metric, we discuss images of these closed timelike curves in the final part of this paper.
@@
In this paper, we focus on the cases of $p=4$ and $p=6$ for the following reasons: first, as illustrated in Fig. 3, as the value of $p$ increases, the parameter range for the QIRK black hole expands, and the orange curve representing the avoidance of CTCs approaches the $a$-axis, with $a_{\text{max}}$ gradually approaching 1. This could potentially obscure certain distinctive features of the QIRK black hole. Second, from the aspect of black hole shadow (not shown here), a smaller value of $p$ results in greater deviations from the Kerr black hole. Thus, considering smaller values of $p$ better highlights the differences between the QIRK black hole and the Kerr black hole.
@@
Table 1: Maximum rotation parameter $a_{\text{max}}$ for different values of $p$. Here $a_{\text{max}}$ refers to the spin at which the $\widetilde{\omega}_{c}$ and $\widetilde{\omega}_{m}$ curves intersect, as shown in Fig. 3.
@@
III Null geodesics
Photons follow null geodesics in QIRK spacetime, and these paths govern lensing and the formation of black hole images. By solving for unstable photon orbits and critical impact parameters, we quantify how quantum corrections alter light deflection relative to Kerr. This is essential for linking the QIRK geometry to observable imaging features.
Before proceeding with the discussion, we would like to clarify the following points: It has been observed that for many rotating black holes, their null geodesic equations share a similar form to that of the Kerr black hole, with the only difference being a correction to $\Delta$. Considering that, first, we have already discussed the behavior of $\Delta$ for the QIRK black hole in Fig. 1, and second, substituting the specific form of $\Delta$ for the QIRK black hole would make the expressions unnecessarily lengthy and less intuitive, we will not include the explicit form of $\Delta$ in this section.